# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v1

@@ 4 Method
Given a corrupted image ${\mathbf{y}}_{0}$, the goal is to sample $\hat{{\mathbf{z}}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ using a pre-trained latent diffusion model as prior. Then, we can transform this sample into an image by applying a pre-trained decoder ${\mathcal{D}}$, i.e. ${\mathbf{x}}_{0}\coloneqq{\mathcal{D}}(\hat{{\mathbf{z}}}_{0})$. We first define a generative model for the data and then apply Bayesian inference on all latent variables using blocked Gibbs sampling and SMC. Specifically, we use the forward diffusion process to augment the model with additional auxiliary observations, and then apply posterior inference using a sequential Monte Carlo (SMC) approach based on the backward diffusion process. The corresponding graphical model can be seen in Figure 3.
@@ 4.1 The Generative Model
We now explicitly define the data generation model based on the forward diffusion process of DDIM (Song et al., 2021a),
$1.~{}{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{0})$
$2.~{}{\mathbf{z}}_{T}|{\mathbf{z}}_{0}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{T}}{\mathbf{z}}_{0},(1-\bar{\alpha}_{T}){\mathbf{I}})),$
$3.~{}{\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0}\sim p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})\quad\forall t\in\{2,...,T\},$
$4.~{}{\mathbf{y}}_{t}|{\mathbf{z}}_{t}\sim\mathcal{N}({\mathcal{A}}(\underbrace{{\mathcal{D}}({\mathbf{z}}_{t}))}_{{\mathbf{x}}_{t}},\tau^{2}{\mathbf{I}})\quad\forall t\in\{0,...,{T-1}\}.$
Here $p({\mathbf{z}}_{0})$ is a prior distribution over ${\mathbf{z}}_{0}$, $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ is defined by the corruption model, and
@@
is defined by the forward diffusion process. Introducing unobserved data is a known technique in statistics for conducting effective Markov chain Monte Carlo (MCMC) sampling (Van Dyk & Meng, 2001; Dou & Song, 2024). In our case, we can use it while leveraging the dependencies between the variables in order to build an efficient SMC sampling procedure, as described in the next section.
@@ 4.2 Sampling Procedure
Given the generative model defined in Section 4.1, our aim is to apply Bayesian inference over the latent variables. In broad strokes, to obtain a sample $\hat{{\mathbf{z}}}_{0}\sim p({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ we use blocked Gibbs sampling to sample in turns from $p({\mathbf{y}}_{1:T-1}|{\mathbf{z}}_{0:T},{\mathbf{y}}_{0})$ and then use SMC to sample from $p({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T-1})$. Specifically, we propose the following procedure:
@@
    Obtain an initial guess for $\hat{{\mathbf{z}}}_{0}$ (detailed in Sec. 4.2.1),
@@
      Sample, ${\mathbf{z}}_{1:T}\sim p({\mathbf{z}}_{1:T}|\hat{{\mathbf{z}}}_{0},{\mathbf{y}}_{0})=p({\mathbf{z}}_{1:T}|\hat{{\mathbf{z}}}_{0})$ according to the forward process of DDIM (Eq. 2).
@@
      Sample, ${\mathbf{y}}_{1:T-1}\sim p({\mathbf{y}}_{1:T-1}|\hat{{\mathbf{z}}}_{0},{\mathbf{z}}_{1:T},{\mathbf{y}}_{0})=p({\mathbf{y}}_{1:T-1}|{\mathbf{z}}_{1:T-1})=\prod_{t=1}^{T-1}\mathcal{N}({\mathbf{y}}_{t}|{\mathcal{A}}({\mathcal{D}}({\mathbf{z}}_{t})),\tau^{2}{\mathbf{I}})$,
    3. (c)
      Sample N particles $\{\hat{{\mathbf{z}}}_{T}^{(i)}\}_{i=1}^{N}\sim\mathcal{N}({\mathbf{z}}_{T},\xi^{2}{\mathbf{I}})$,
    4. (d)
      Sample $\hat{{\mathbf{z}}}_{0:T-1}\sim p_{\theta}({\mathbf{z}}_{0:T-1}|{\mathbf{y}}_{0:T-1},\{\hat{{\mathbf{z}}}_{T}^{(i)}\}_{i=1}^{N})$ using SMC based on a pre-trained diffusion model and select one particle $\hat{{\mathbf{z}}}_{0}$ for the next iteration (detailed in Sec. 4.2.2).
Here we use the dynamics of the forward process and the graphical model dependencies in steps (a) and (b). In step (c) we sample from a Gaussian centered at the final ${\mathbf{z}}_{T}$ having a small std $\xi$ instead of the prior distribution as commonly done and described in Section 2. Although empirically we found that this step did not have a significant effect on the final model performance, we did it to be better aligned with the generated ${\mathbf{y}}_{t}$ from the forward process. In addition, we found that one forward-backward pass suffices to achieve good results. Hence, in our empirical evaluations, we apply Step 2 only once. In conclusion, steps (a), (b) and (c) are fully defined. The two steps that are not straightforward are how to obtain an initial guess for $\hat{{\mathbf{z}}}_{0}$ (step 1) and how to perform the sampling process in step (d). We discuss both next.
4.2.1 Initial Guess for $\hat{{\mathbf{z}}}_{0}$
The first challenge is to obtain an initial $\hat{{\mathbf{z}}}_{0}$ (step 1. in the sampling procedure). To reduce the variance in the process and accelerate convergence, we performed the following optimization procedure in pixel space:
$\hat{{\mathbf{x}}}_{0}$ $=\operatorname*{arg\,max}_{\mathbf{x}}log~{}p_{\theta}({\mathbf{y}}_{0}|{\mathbf{x}})=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2},$ (3)
and then we applied the encoder on the outcome, namely $\hat{{\mathbf{z}}}_{0}={\mathcal{E}}(\hat{{\mathbf{x}}}_{0})$. For linear inverse problems, this optimization problem can be solved in closed form (Song et al., 2024; Wang et al., 2023), although it can be costly, as it requires inverting the linear operator ${\mathcal{A}}$. An alternative for this procedure is to apply the optimization process directly in the latent space. However, in our experiments we found that the former option worked better and it did not involve expensive gradient propagation through the decoder.
@@ 4.2.2 Posterior Sampling
We now move on to explain step (d) of the sampling procedure. Due to the non-linearity of the decoder, even for linear inverse problems, finding the exact posterior is intractable. One option to overcome this difficulty is to use SMC sampling. In what follows, we first describe an approximate posterior for $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T-1})$, then we suggest an iterative procedure based on SMC sampling.
@@
First, notice that because of the structure of the model, the posterior density of the r.v. ${\mathbf{z}}_{t}$ at each step $t$ depends only on ${\mathbf{z}}_{t+1:T}$. Hence, only $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T-1})$ needs to be computed at each time step $t$. However, even computing an unnormalized quantity of that posterior can be costly. Therefore, we make the following assumption $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T-1})\approx p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T-1},{\mathbf{y}}_{0})$. We assume this as ${\mathbf{y}}_{0}$ stores all the input information. Now we can arrive at the following recursive formula (to prevent cluttered notations we omit here the subscript of $\theta$ from the probability densities):
$p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T},{\mathbf{y}}_{0})$
$\propto p({\mathbf{y}}_{t}|{\mathbf{z}}_{t:T},{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})p({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1:T},{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1},{\mathbf{y}}_{0})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$=p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})\frac{p({\mathbf{y}}_{0}|{\mathbf{z}}_{t})}{p({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})}p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0})$
$\approx\frac{p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))}{p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t+1}))}p({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0}).$
Where, in the first transition we used Bayes rule, in the third transition we used the Markovian assumption, in the forth transition we used Bayes rule again, and in the last transition we make an additional approximation and condition on the posterior mean estimator for both time $t$ and time $t+1$. We define $p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))=\mathcal{N}({\mathbf{y}}_{0}|{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))),(1-\bar{\alpha}_{t}){\mathbf{I}}))$ and similarly for $p({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t+1}))$, where as in (Wu et al., 2024) the variance term is taken to be the variance of the forward diffusion process.
Figure 4: Comparison between LD-SMC and baseline methods on Gaussian deblurring of images from the FFHQ dataset.
Now, we can derive an SMC procedure using the proposed posterior. Specifically, we used sequential importance sampling with resampling (SISR). The general prescription (adapted to our setting) is as follows (Särkkä, 2013):
  1. 1.
    Collect N particles as detailed in step (c) in Section 4.2.
  2. 2.
    For $t=T-1,...,0$:
    1. i.
      Draw $\{\tilde{{\mathbf{z}}}_{t}^{(i)}\}_{i=1}^{N}$ samples from a proposal distribution (detailed in Sec. 4.2.3):
      $\tilde{{\mathbf{z}}}_{t}^{(i)}\sim\pi(\tilde{{\mathbf{z}}}_{t}|\hat{{\mathbf{z}}}_{t+1}^{(i)},{\mathbf{y}}_{0:T-1})\quad i=1,...,N.$
    2. ii.
      Compute the proposal weights for all $i\in\{1,...,N\}$ according to:
      $w_{t}^{(i)}$ $\propto\frac{p_{\theta}({\mathbf{y}}_{t}|\tilde{{\mathbf{z}}}_{t}^{(i)})p_{\theta}({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}(\tilde{{\mathbf{z}}}_{t}^{(i)}))p_{\theta}(\tilde{{\mathbf{z}}}_{t}^{(i)}|\hat{{\mathbf{z}}}_{t+1}^{(i)})}{\pi(\tilde{{\mathbf{z}}}^{(i)}_{t}|\hat{{\mathbf{z}}}_{t+1}^{(i)},{\mathbf{y}}_{0:T-1})p_{\theta}({\mathbf{y}}_{t}|\bar{{\mathbf{z}}}_{0}(\hat{{\mathbf{z}}}_{t+1}^{(i)}))},$
      and then normalize the weights to sum to one.
    3. iii.
      Resample $N$ particles (with replacements) from the discrete distribution: $\hat{{\mathbf{z}}}_{t}^{(i)}\sim\{\tilde{{\mathbf{z}}}_{t}^{(1)},...,\tilde{{\mathbf{z}}}_{t}^{(N)}\}$ with weights $\{w_{t}^{(1)},...,w_{t}^{(N)}\}$.
Here, $p_{\theta}(\tilde{{\mathbf{z}}}_{t}^{(i)}|\hat{{\mathbf{z}}}_{t+1}^{(i)})$ is defined according to Eq. 1 and we assume that the resampling step is performed at each iteration. While the derivation is different, the resulting weighting scheme of LD-SMC bears some resemblance to the weighting of TDS (Wu et al., 2024). The main difference between the two methods is the dependence on the auxiliary variables ${\mathbf{y}}_{1:T}$ in the proposal distribution and resampling weights. Empirically, we observed that this additional conditioning helped to better align the sampling with the corrupted image ${\mathbf{y}}_{0}$ compared to using the posterior mean approximation as in (Chung et al., 2023b) and (Wu et al., 2024). Finally, to select one sample at time $t=0$ for terminating the algorithm in Section 4.2 or the next iteration in the Gibbs sampling procedure, we take the particle that has the highest weight.

### v2

@@ 2 Background
3 Related Work
Inverse problems have a long and evolving history, with methodologies that have undergone significant advances over the years (Daras et al., 2024). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021d) have emerged as effective priors for solving inverse problems in image data (Choi et al., 2021; Kawar et al., 2022; Chung et al., 2023a, b; Rout et al., 2023; Song et al., 2023; Wang et al., 2023; Dou & Song, 2024; Garber & Tirer, 2024; Mardani et al., 2024; Sun et al., 2024).
In (Song et al., 2021d) it was shown that to sample from the posterior distribution, $p({\mathbf{x}}_{0}|{\mathbf{y}})$, one can solve a stochastic differential equation based on the prior score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{x}}_{t})$, and the conditional score, $\nabla_{{\mathbf{x}}_{t}}\log~{}p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})$. While the first term is easy to compute, the latter term requires integration over the full diffusion path from time $t$ to 00. A useful and easy-to-calculate approximation found in several studies is $p_{\theta}({\mathbf{y}}|{\mathbf{x}}_{t})\approx p_{\theta}({\mathbf{y}}|\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}])\approx p_{\theta}({\mathbf{y}}|\bar{{\mathbf{x}}}_{0}({\mathbf{x}}_{t}))$, which is readily available at each step (Chung et al., 2023b; Song et al., 2023; Wu et al., 2024). Specifically, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b) uses this approximation for linear and nonlinear inverse problems with Gaussian and Poisson likelihood models. Other methods also utilize the pseudoinverse of the corruption operator ${\mathcal{A}}$ (Tirer & Giryes, 2018). $\Pi$GDM (Song et al., 2023) introduces a guidance scheme by matching the denoised output and the corrupted image ${\mathbf{y}}$, via a transformation of both. DDNM (Wang et al., 2023) suggested refining the contents of the null space of ${\mathcal{A}}$ during the backward diffusion process. A different approach proposed in (Mardani et al., 2024) is to approximate the posterior using a variational approach based on the score matching objective presented in (Song et al., 2021b). An additional category of inverse problem approaches that use diffusion models is designed with the objective of asymptotic exactness (Cardoso et al., 2023; Trippe et al., 2023; Wu et al., 2024; Dou & Song, 2024). These methods utilize SMC techniques, which have also been applied for unconditional sampling (Chen et al., 2025), to target exact sampling from the posterior distribution $p({\mathbf{x}}_{0}|{\mathbf{y}})$. SMC-Diff (Trippe et al., 2023) was designed mainly for motif-scaffolding. It uses the prior distribution as a proposal which is known to require a large number of particles for accurate estimation results (Särkkä, 2013), a severe limitation for latent space diffusion models due to expensive decoder evaluations. We experimented with this proposal in our initial attempts and witnessed that even $150$ particles were not enough to sample plausible reconstructions. MCGDiff (Cardoso et al., 2023) was designed for linear inverse problems only, which makes it unsuitable for inverse problems with latent-space diffusion models due to encoder-decoder involvement. TDS (Wu et al., 2024), a recent SMC-based method, solves general inverse problem tasks using the twisting technique. This method also uses the approximation of DPS, but by applying SMC sampling it can correct for it. FPS (Dou & Song, 2024), also a recent SMC-based method, uses auxiliary variables. FPS generates a sequence of observations ${\mathbf{y}}_{1:T}$ based on a duplex diffusion process, one process at the ${\mathbf{x}}$ space and the other process at the ${\mathbf{y}}$ space. Since this method is designed for linear inverse problems only, it permits tractable Bayesian inference. Our method combines the ideas of both TDS and FPS to obtain the best of both. Namely, we use the posterior mean approximation and ${\mathbf{y}}_{1:T}$ in our SMC sampling process. As we will show, this combination can be helpful in both understanding the general semantics of an image and capturing fine details. We note here that there are other methods that applied label augmentations similar to FPS, one example is (Abu-Hussein et al., 2022) which is applicable for linear inverse problems only. Hence, this method as well cannot be trivially combined with latent space diffusion models.
Algorithm 1 LD-SMC
1: Set ${\mathbf{z}}_{0}={\mathcal{E}}({\mathbf{x}}_{0})$, where ${\mathbf{x}}_{0}=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2}$
2: Sample ${\mathbf{z}}_{1:T}\sim p({\mathbf{z}}_{1:T}|{\mathbf{z}}_{0})$ according to the forward
    process of DDIM (Eq. 2)
3: For $k=1,...,K$:
4:  Sample ${\mathbf{y}}_{1:T}\sim\prod_{t=1}^{T}\mathcal{N}({\mathbf{y}}_{t}|{\mathcal{A}}({\mathcal{D}}({\mathbf{z}}_{t})),\tau^{2}{\mathbf{I}})$
5:  Sample ${\mathbf{z}}_{0:T}\sim p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ using $\mathbf{SISR}({\mathbf{y}}_{0:T})$
6: $\mathbf{SISR}({\mathbf{y}}_{0:T})$:
7:    For $i=1,...,N$:
8:   Sample ${\mathbf{z}}_{T}^{(i)}\sim p_{\theta}({\mathbf{z}}_{T})$ and initialize ${\mathbf{z}}^{(i)}=({\mathbf{z}}_{T}^{(i)})$
9:   Set $w_{T}^{(i)}$ according to Eq. 6
10:  Normalize $\{w_{T}^{(i)}\}_{i=1}^{N}$ to sum to one
11:  For $t=T-1,...,0$:
12:   $\{{\mathbf{z}}^{(i)}\}_{i=1}^{N}\sim\text{Multi}(\{{\mathbf{z}}^{(i)}\}_{i=1}^{N},\{w_{t+1}^{(i)}\}_{i=1}^{N})$,
            a resampling step
13:   For $i=1,...,N$:
14:    Sample ${\mathbf{z}}_{t}^{(i)}\sim\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1}^{(i)})$ (Eq. 7) # proposal
15:    Set $w_{t}^{(i)}$ according to Eq. 6
16:    Update ${\mathbf{z}}^{(i)}=({\mathbf{z}}_{t}^{(i)},...,{\mathbf{z}}_{T}^{(i)})$
17:   Normalize $\{w_{t}^{(i)}\}_{i=1}^{N}$ to sum to one
18:  Sample one chain ${\mathbf{z}}\sim\text{Multi}(\{{\mathbf{z}}^{(i)}\}_{i=1}^{N},\{w_{0}^{(i)}\}_{i=1}^{N})$
19:  Return ${\mathbf{z}}$
Figure 3: Comparison between LD-SMC and baseline methods on inpainting of ImageNet images.
Several inverse sampling methods were specifically tailored for latent diffusion models. PSLD (Rout et al., 2023) extend DPS (Chung et al., 2023b) by incorporating an additional gradient update step to guide the diffusion process to sample latent representations that maintain the integrity of the decoding-encoding transformation, ensuring it remains non-lossy. STSL (Rout et al., 2024) uses second-order information for reconstruction based on efficient approximations. Comparative analysis with STSL was not feasible due to the absence of publicly available code, making replication challenging. ReSample (Song et al., 2024), a contemporary method alongside PSLD, acts in two stages; it first applies hard data consistency to obtain latent variables that are consistent with the observed measurements and then employs a resampling scheme to map the sample back onto the data manifold. Concurrent to this study Nazemi et al. (2024) proposed a particle filtering approach. Their method builds on PSLD and DPS update in the proposal distribution. Similarly to TDS (Wu et al., 2024) the connection to the labels is only through ${\mathbf{z}}_{0}$ using the approximate mean estimator. Also, unlike LD-SMC it does not enjoy asymptotic guarantees. Recently, both DAPS (Zhang et al., 2024) and MGPS (Moufad et al., 2024) have been proposed to address inverse problems both in pixel and latent space. DAPS decouples latents in the sampling trajectory by first sampling from ${\mathbf{z}}_{0}|{\mathbf{z}}_{t+\Delta t},{\mathbf{y}}$ using MCMC techniques, for some $\Delta t>0$, and then sampling ${\mathbf{z}}_{t}$. MGPS also decomposes the sampling, but to an intermediate midpoint and uses variational inference for posterior sampling. Given the similarities of both methods, in our experiments we compare only to the former. SILO (Raphaeli et al., 2025), an additional recent study, first learns a NN that mimics an encoder operation on labels ${\mathbf{y}}$ which is then used to guide the sampling process. In our empirical evaluation, we compare to methods that apply sampling only.
@@ 4 Method
Given a corrupted image ${\mathbf{y}}_{0}$, the goal is to sample ${\mathbf{z}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ using a pre-trained latent diffusion model as prior. Then, we can transform this sample into an image by applying a pre-trained decoder ${\mathcal{D}}$, i.e. ${\mathbf{x}}_{0}\coloneqq{\mathcal{D}}({\mathbf{z}}_{0})$. Here, we stack all the parameters of the networks (diffusion and encoder-decoder) under $\theta$. In what follows, we first define a generative model for the data and then apply Bayesian inference on all latent variables using blocked Gibbs sampling and SMC. Specifically, we use the diffusion process to augment the model with additional auxiliary observations and apply posterior inference using a sequential Monte Carlo (SMC) approach based on the backward diffusion process. The corresponding graphical model can be seen in Figure 2.
@@ 4.1 The Generative Model
We now explicitly define the data generation model,
$1.~{}{\mathbf{z}}_{T}\sim\mathcal{N}(0,{\mathbf{I}})),$
$2.~{}{\mathbf{z}}_{t-1}|{\mathbf{z}}_{t}\sim p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})\quad\forall t\in\{1,...,T\},$
$3.~{}{\mathbf{y}}_{t}|{\mathbf{z}}_{t}\sim\mathcal{N}({\mathcal{A}}(\underbrace{{\mathcal{D}}({\mathbf{z}}_{t})}_{{\mathbf{x}}_{t}}),\tau^{2}{\mathbf{I}})\quad\forall t\in\{0,...,{T}\}.$
Where, $p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ is defined by the corruption model, $p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t})=\mathbb{E}_{p({\mathbf{z}}_{0}|{\mathbf{z}}_{t})}[p({\mathbf{z}}_{t-1}|{\mathbf{z}}_{t},{\mathbf{z}}_{0})]$ is a backwards generative process that corresponds to a non-Markovian forward diffusion process with
@@
set according DDIM (Song et al., 2021a). Introducing unobserved data is a known technique in statistics for conducting effective Markov chain Monte Carlo (MCMC) sampling (Van Dyk & Meng, 2001; Dou & Song, 2024). In our case, we can use it while leveraging the dependencies between the variables in order to build an efficient SMC sampling procedure, as described in the next section.
@@ 4.2 Sampling Procedure
Given the generative model defined in Section 4.1, our aim is to apply Bayesian inference over the latent variables. In broad strokes, to obtain a sample ${\mathbf{z}}_{0}\sim p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ we use blocked Gibbs sampling to sample in turns from $p_{\theta}({\mathbf{y}}_{1:T}|{\mathbf{z}}_{0:T},{\mathbf{y}}_{0})$ using knowledge on the corruption model and $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ using SMC. Specifically, we propose the following procedure:
@@
    Obtain an initial guess for ${\mathbf{z}}_{0}$ (Sec. 4.2.1).
@@
    Sample, ${\mathbf{z}}_{1:T}\sim p({\mathbf{z}}_{1:T}|{\mathbf{z}}_{0},{\mathbf{y}}_{0})=p({\mathbf{z}}_{1:T}|{\mathbf{z}}_{0})$ according to the forward process of DDIM (Eq. 2).
  3. 3.
@@
      Sample, ${\mathbf{y}}_{1:T}\sim p_{\theta}({\mathbf{y}}_{1:T}|{\mathbf{z}}_{0:T},{\mathbf{y}}_{0})=p_{\theta}({\mathbf{y}}_{1:T}|{\mathbf{z}}_{1:T})=\prod_{t=1}^{T}\mathcal{N}({\mathbf{y}}_{t}|{\mathcal{A}}({\mathcal{D}}({\mathbf{z}}_{t})),\tau^{2}{\mathbf{I}})$.
@@
      Sample, ${\mathbf{z}}_{0:T}\sim p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ using SMC based on a pre-trained diffusion model (Sec. 4.2.2).
Here, we use the dynamics of the forward process and the graphical model dependencies in steps 2 and 3(a). The two steps that are not straightforward are obtaining an initial guess for ${\mathbf{z}}_{0}$ (step 1) and performing the sampling process in step 3(b). We discuss both in the following sections, but before that we make a few comments. First, clearly the run-time of the algorithm depends linearly on the number of Gibbs iterations. As a result, the sampling time can be slow. However, with proper initialization (as discussed in Section 4.2.1), we empirically found that even one iteration of Step 3 suffices to achieve good results. Hence, unless otherwise stated, in our empirical evaluations, we applied this step only once. Second, one may be concerned that evaluating the decoder on noisy latent variables can generate non-natural images. While this may be true, empirically we observed a graceful degradation in image quality with time, as seen in Figure 7. Furthermore, the latent variables, even if noisy, carry information about ${\mathbf{y}}_{0}$. This information is then transferred to the auxiliary labels which help guide the sampling process.
Figure 4: Comparison between LD-SMC and baseline methods on Gaussian deblurring of ImageNet images.
4.2.1 Initial Guess for ${\mathbf{z}}_{0}$
The first challenge is to obtain an initial ${\mathbf{z}}_{0}$ (step 1. in the sampling procedure). There are multiple sensible ways to perform this initialization. Importantly, ${\mathbf{z}}_{0}$ should carry information about the measurement $\mathbf{y}_{0}$, which can then be used to guide the sampling process using the auxiliary labels. As a result, the variance in the process will be reduced and the convergence will be accelerated. For example, one option is to leverage the pseudoinverse of ${\mathcal{A}}$ for linear operators similar to (Wang et al., 2023). But, that would not work for nonlinear corruption operators. Although there are mitigations for this problem (Song et al., 2023), to maintain generality, we applied the following optimization procedure,
${\mathbf{x}}_{0}$ $=\operatorname*{arg\,max}_{\mathbf{x}}log~{}p_{\theta}({\mathbf{y}}_{0}|{\mathbf{x}})=\operatorname*{arg\,min}_{\mathbf{x}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathbf{x}})||^{2}_{2},$ (3)
where ${\mathbf{x}}$ is initialized to ${\mathcal{D}}(\hat{{\mathbf{z}}}_{0})$ for $\hat{{\mathbf{z}}}_{0}\sim\mathcal{N}(0,{\mathbf{I}})$. Then, after reaching convergence, an initial ${\mathbf{z}}_{0}$ is obtained by applying the encoder on the outcome, namely ${\mathbf{z}}_{0}={\mathcal{E}}({\mathbf{x}}_{0})$. An alternative for this procedure is to perform the optimization process directly in the latent space. However, in our experiments we found that the former option worked better and was substantially faster as it did not involve expensive gradient propagation through the decoder.
@@ 4.2.2 Posterior Sampling
We now move on to explain step 3(b) of the sampling procedure. Due to the non-linearity of the decoder, even for linear corruption operators ${\mathcal{A}}$, finding the exact posterior is intractable. One option to overcome this difficulty is to use SMC sampling. In what follows, we first describe approximate (tractable) target distributions for $\{p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T})\}_{t=0}^{t=T}$, then we present the induced sequence weights $w_{t}$ and a novel proposal distributions $\pi_{t}$ for all timesteps $t$. These components are used as part of a general sequential importance sampling with resampling (SISR) algorithm as depicted in Algorithm 1.
@@
Approximate target distributions. First, notice that due to the structure of the model, the posterior density of the r.v. ${\mathbf{z}}_{t}$ at each step $t$ depends only on ${\mathbf{z}}_{t+1:T}$. Hence, the target distribution of interest at each timestep $t$ is $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T})$. However, even computing the unnormalized quantity of that target distribution can be costly. Therefore, we make the following assumption $p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{0:T})\approx p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T},{\mathbf{y}}_{0})$. This assumption is reasonable since ${\mathbf{y}}_{0}$ stores all the input information to begin with. In Appendix B we arrive at the following recursive formula which forms the approximate target distributions of the SMC procedure for all $t$:
$p_{\theta}({\mathbf{z}}_{t:T}|{\mathbf{y}}_{t:T},{\mathbf{y}}_{0})\propto$ (4)
$\begin{cases}p_{\theta}({\mathbf{y}}_{T}|{\mathbf{z}}_{T})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{T})p({\mathbf{z}}_{T}),\qquad\qquad\qquad\qquad t=T,\\ \frac{p_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t})\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})}p_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})p_{\theta}({\mathbf{z}}_{t+1:T}|{\mathbf{y}}_{t+1:T},{\mathbf{y}}_{0}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad~{}~{}0<t<T,\\ \frac{p_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})}{\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{1})}p_{\theta}({\mathbf{z}}_{0}|{\mathbf{z}}_{1})p_{\theta}({\mathbf{z}}_{1:T}|{\mathbf{y}}_{0:T}),\qquad\qquad\qquad t=0,\end{cases}$
where we define $\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t})\coloneqq p_{\theta}({\mathbf{y}}_{0}|\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))=\mathcal{N}({\mathbf{y}}_{0}|{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t}))),(1-\bar{\alpha}_{t}){\mathbf{I}}))$, with the variance taken to be the variance of ${\mathbf{z}}_{t}|{\mathbf{z}}_{0}$ according to the forward diffusion process. Importantly, as the final target distribution at $t=0$ matches the desired distribution $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ (see Eq. 10 in Appendix C), in the large compute limit samples from the correct target can be obtained despite all approximations.